Probl below. em 7. The annual tuitions of 16 randomly selected 2-year public colleges are described by the boxplot Tutions a. Is the distribution of tuitions symmetric, skewed left, or skewed right? (1 pt.) Lutf sKwed b. Estimate the percentage of the college tuitions cost between 24 and 28 thousand. (2 pts.) 20 30 Tultions (in thousand dollars) 추@s_Qィ 2 c. How many colleges have tuitions cost between 24 and 28 thousand? (2 pts.)

Need help ?!
Answer #1

Ans:

b)From box plot:

Q1=24

Q3=29

IQR=29-24=5

Between 24 and 29, there will be 50% of the data.

So,between 24 and 28,there will be approximately 40% of the data.

hence, 40% of the college tuitions cost between 24 and 28 thousand.

c)

Number of college having tuitions between 24 and 28=0.4*16=6.4